OpenStudy (anonymous):

Solve: |x + 3| = 3 Answer {0, 6} {0, -6} {-6, 6} {-6}

OpenStudy (jackellyn):

You have to solve this by putting the 3 as negative and positive. Here is an example: |x-4|=6 x-4=6 x-4=-6 x=10 x=-2 Get it?

OpenStudy (anonymous):

No.

OpenStudy (jackellyn):

You have to do two cases when you have absolute values. Case 1 look just like the original problem but without the absolute value. x+3=3 x=3-3=? Case 2 is where the number on the right changes negative x+3=-3 x=-3-3=?